the number of silver atoms present in a 90% pure silver wire weighing 10gm is

the number of  silver atoms present in a 90% pure silver  wire weighing 10gm is 

Grade:11

1 Answers

somi teez
6 years ago
hi,
  atomic mass of silver=108 g
 amount of silver present in 10 gm=90% of 10 gm =9 gm
no. of moles=9/108=1/12
in 1 mole , 6.023 *10^23 atomse present
 hence, in 1/12 moles no of atoms present=1/12*6.023*10^23....answer